Virginia Company, a merchandising firm, operated 4 sales offices last year at a total cost of $560,000, of which $76,000 represented fixed costs. Virginia has determined that total costs are significantly influenced by the number of sales offices operated. Last year's costs and number of sales offices can be used as the basis for predicting annual costs. What would be the budgeted cost for the coming year if Virginia were to operate 7 sales offices? (CPA adapted)

Multiple Choice

  • $847,000.

  • $923,000.

  • $802,000.

  • $608,000.
Answer #1
Variable cost for 4 sales office 484000 =560000-76000
Variable cost per sales office 121000 =484000/4
Variable cost for 7 sales office 847000 =121000*7
Add: Fixed costs 76000
Budgeted cost for the coming year 923000
Option B $923,000 is correct